How many cups of grapefruit juice must be added to 30 cups of a punch that is 8% grapefruit juice to obtain a punch that is 10% grapefruit juice?

Answer by stanbon(75887)

How many cups of grapefruit juice must be added to 30 cups of a punch that is 8% grapefruit juice to obtain a punch that is 10% grapefruit juice?
-------------
Equation:
grape + grape = grape
----
x + 0.08*30 = 0.10(x+30)
---
Multiply thru by 100 to get:
100x + 8*30 = 10x + 10*30
90x = 2*30
x = 2/3 of a cup of 100% grapejuice must be added.
======================
Cheers,
Stan H.